A 1030 kg car is pulling a 290 kg trailer. Together the car and trailer move forward with an acceleration of 2.13 m/s2. Ignore any frictional force of air drag on the car and all frictional forces on the trailer.

(a) Determine the net force on the car.

Magnitude:

Direction: Forward/Backward

b.)Determine the net force on the trailer:

Magnitude

Direction: Forward/Backward

c.) Determine the force exerted by the trailer on the car.

Magnitude:

Direction: Forward/Backward

d.) Determine the resultant force exerted by the car on the road. (Assume that the forward direction is along the +x-direction.)

Magnitude:

Direction: Forward/Backward

*****Please Post step by step guide, written on paper preferred but whichever way you like is ok too******
Answer #1

a) F=M*a =1030*2.13 =2193.9 N Forward
b) F=m*a= 617.7 N Forward
c)F=617.7 N backward
d) F= (M+m)*a = 2811.6 N Backward